sum of 3 consecutive natural numbers each divisible by 3 is 72 .what is largest among them?
Answer by richwmiller(17219)

x+x+2+x+4=72
3x+6=72
3x=66
x=22
x+2=24
x+4=26
"Each divisible by 3 " makes no sense!